OpenStudy (daisyduck04):

help fan and medal :fractions 4/9 divided by 3/7

jabez177 (jabez177):

\[\frac{ 4 }{ 9 }\times \frac{ 7 }{ 3 }\]

jabez177 (jabez177):

That is another way to do it. What is 4 x 7?

OpenStudy (daisyduck04):

28

jabez177 (jabez177):

Now what is 9 x 3?

OpenStudy (daisyduck04):

27

OpenStudy (daisyduck04):

oh thx

jabez177 (jabez177):